OpenStudy (anonymous):

Using synthetic division, what is the quotient: (3x3 + 4x - 32) ÷ (x - 2)? 3x2 + 6x - 16 3x2 - 6x + 16 3x2 + 6x + 16 3x2 - 6x - 16 + 1 over the quantity of x minus 2

OpenStudy (amistre64):

can you show your work?

OpenStudy (anonymous):

ok \[3x ^{3}\div x=2x ^{2}\]\[32\div2=16\]

OpenStudy (anonymous):

\[4x \div x= 3x\]

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

that kinda looks like you are attempting a long hand division .... which is fine to get an option with

OpenStudy (amistre64):

my strategy is this: (3x3 + 4x - 32) ÷ (x - 2) ^^^^ this equals 0 when x=2 now make sure ALL the "x" parts are present 3 2 1 0 3x3 + 4x - 32 we are missing an x^2 part, fill it in with 0x^2 3 2 1 0 3x3 +0x2 + 4x - 32 now we can run thru the synthetic method

OpenStudy (anonymous):

ok i follow so far

OpenStudy (amistre64):

i use 3 rows like this 3x3 +0x2 + 4x - 32 0 <-- add this row down -------------------- 2 ) <-- multiply this row to move across, using the 2 as the multiplier

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

3x3 +0x2 + 4x - 32 0 -------------------- 2 ) 3 3*2 = 6, so plug it in and add down 3x3 +0x2 + 4x - 32 0 6 -------------------- 2 ) 3 6 16 2*6 = 12, so plug it in and add down 3x3 +0x2 + 4x - 32 0 6 12 -------------------- 2 ) 3 6 16 what would the next step be? its very repetitive

OpenStudy (anonymous):

you would plug 16 in and add it to 32?

OpenStudy (amistre64):

2 times 16, then add that result to -32

OpenStudy (anonymous):

ok so it would get 0 on the end

OpenStudy (amistre64):

yes, now we can see the parts: 3x3 +0x2 + 4x - 32 0 6 12 16 -------------------- 2 ) 3 6 16 0 ^______________^ ^remainder quotient and this is why i keep the x parts in place, they remind me how to rework the quotient ... 3x3 +0x2 + 4x - 32 0 6 12 16 -------------------- 2 ) 3 6 16 0 x^2 x^1 x^0 3x^2 +6x +16

OpenStudy (anonymous):

thank you so much man i have been stuck on this for a day

OpenStudy (amistre64):

youre welcome :)